# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v1

@@ 1.2 Contributions
We leverage recent results from distributionally robust optimization and convex learning to propose a new Wasserstein distributionally robust shallow convex neural network (WaDiRo-SCNN) with out-of-sample performance guarantees and simple low-stochasticity training. The training is formulated as a convex optimization problem efficiently solvable with open-source solvers. This training problem can be constrained to account for hard physical constraints as long as they are convex. We first benchmark our model with a numerical study in a controlled synthetic environment. We then apply our approach to distribution feeder-specific impact prediction of non-residential buildings enrolled in a demand response program. To the best of our knowledge, our specific contributions are as follows:
  • We demonstrate that the SCNN with $\ell_{1}$-loss training problem is similar to a linear regression over a modified sample space.
  • We formulate an exact, tractable, and low-stochasticity Wasserstein distributionally robust training program for shallow convex neural networks under both $\ell_{1}$-norm and $\ell_{2}$-norm order-1 Wasserstein distances.
  • We derive out-of-sample guarantees for WaDiRo-SCNNs with $\ell_{1}$-loss training via Rademacher complexity.
@@ 2.1 Distributionally Robust Optimization
Distributionally robust optimization (DRO) minimizes a worst-case expected loss function over an ambiguity set, i.e., an uncertainty set in the space of probability distributions [21]. The problem can be formulated as:
$\inf_{\bm{\beta}}\sup_{\mathbb{Q}\in\Xi}\ \mathbb{E}^{\mathbb{Q}}[h_{\bm{\beta}}(\mathbf{z})],$ (DRO)
where $\mathbf{z}\in\mathcal{Z}\subseteq\mathbb{R}^{d},d\in\mathbb{N}$, is a random sample, $\mathcal{Z}$ is the set of all possible values of $\mathbf{z}$, $\mathbb{Q}$ is a distribution from the ambiguity set $\Xi$, $\bm{\beta}\in\mathbb{R}^{p}$ is the $p$-dimension decision vector, and $h_{\bm{\beta}}(\mathbf{z}):\mathcal{Z}\times\mathbb{R}^{p}\mapsto\mathbb{R}$ is the loss function when applying $\bm{\beta}$ on a sample [36, 25].
We are interested in a data-driven branch of DRO which defines the ambiguity set as a Wasserstein ball centered on an empirical distribution containing $N$ samples: ^$\hat{\mathbb{P}}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathbf{z}_{i}}(\mathbf{z})$, where $\delta_{\mathbf{z}_{i}}(\cdot)$ is a Dirac delta function assigning a probability mass of one at each known sample $\mathbf{z}_{i}$. Note that as the number of samples goes to infinity the empirical distribution converges to the true distribution $\mathbb{P}$. The Wasserstein distance, also called optimal transport, can be pictured as the minimal effort that it would take to displace a given pile ($d_{1}$) of a weighted resource, e.g., dirt, to shape a second specific pile ($d_{2}$). The effort of a trip is given by the distance travelled and the weight of the carried resource. Specifically, it provides a sense of similarity between the two distributions. Formally, the order-$t$ Wasserstein distance between distributions $\mathbb{U}$ and $\mathbb{V}$ is:
$W_{\|\cdot\|,t}(\mathbb{U},\mathbb{V})=\left(\min_{\pi\in\mathcal{P}(\mathcal{Z}\times\mathcal{Z})}\int_{\mathcal{Z}\times\mathcal{Z}}\|\mathbf{z}_{1}-\mathbf{z}_{2}\|^{t}\text{d}\pi(\mathbf{z}_{1},\mathbf{z}_{2})\right)^{1/t},$
where $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$ are the marginals of $\mathbb{U}$ and $\mathbb{V}$, respectively [36].
@@
where $\Omega$ is the order-$t$ Wasserstein sphere of radius $\epsilon>0$ centered on ^$\hat{\mathbb{P}}_{N}$, and $\mathcal{P}(\mathcal{Z})$ is the space of all probability distributions with support on $\mathcal{Z}$. In a way, $\epsilon$ can be interpreted as a hyperparameter that controls our conservatism towards ^$\hat{\mathbb{P}}_{N}$. This problem is not tractable [25].
Using the strong duality theorem and the order-1 Wasserstein metric, [25] showed that a tractable reformulation of (WDRO) is obtainable if $h_{\bm{\beta}}(\mathbf{z})$ is convex in $\mathbf{z}\in\mathcal{Z}=\mathbb{R}^{d}$, where $d$ is the dimension of the samples. The problem to consider then becomes:
@@
where $\|\cdot\|_{*}$ is the dual norm $\|\bm{\theta}\|_{*}\equiv\sup_{\|\mathbf{z}\|\leq 1}\bm{\theta}^{\top}\mathbf{z}$ and $h_{\bm{\beta}}^{*}(\bm{\theta})$ is the convex conjugate of $h_{\bm{\beta}}(\mathbf{z})$, $h_{\bm{\beta}}^{*}(\bm{\theta})=\sup_{\mathbf{z}}\{\bm{\theta}^{\top}\mathbf{z}-h_{\bm{\beta}}(\mathbf{z})\}$ [37].
Training machine learning models with (WDRO) has been proved to be a generalization of regularization [38]. By training over the worst-case expected distribution in a Wasserstein ball centered on the training dataset, we avoid overfitting on the empirical distribution and become more robust against outliers, anomalies, and some adversarial attacks.
@@ 2.2 Shallow Convex Neural Networks
Let $\mathbf{X}\in\mathbb{R}^{N\times d}$ be a data matrix of $N$ feature vectors, $\mathbf{x}_{i}\in\mathbb{R}^{d}\ \forall i\in\{1,2,\ldots,N\}\equiv\llbracket N\rrbracket$, and $\mathbf{y}\in\mathbb{R}^{N}$ be the corresponding one-dimensional labels. Define data samples as a feature-label pair: $\mathbf{z}_{i}=(\mathbf{x}_{i},y_{i})\ \forall i\in\llbracket N\rrbracket$. The non-convex training problem of a single-output feedforward shallow neural network (SNN) with $\operatorname{ReLU}$ activation functions is as follows:
$p^{\star}=\min_{\mathbf{W}_{1},\mathbf{w}_{2}}\mathcal{L}\left(\mathbf{y},\sum_{i=1}^{m}\{(\mathbf{X}\mathbf{W}_{1i})_{+}w_{2i}\}\right),$ (SNNT)
where $m$ is the number of hidden neurons, $\mathbf{W}_{1}\in\mathbb{R}^{d\times m}$, and $\mathbf{w}_{2}\in\mathbb{R}^{m}$ are the weights of the first and second layers, $\mathcal{L}:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is a convex loss function, and $(\cdot)_{+}$ is an element-wise $\max\{0,\cdot\}$ operator.
The convex reformulation of the SNN training problem (SNNT) is based on an enumeration process of all the possible $\operatorname{ReLU}$ activation patterns in the hidden layer for a fixed training dataset $\mathbf{X}$ [18]. The enumeration is equivalent to generating all possible hyperplane arrangements passing through the origin and clustering the data with them. The set of all possible activation patterns a $\operatorname{ReLU}$ NN can take for a training set $\mathbf{X}$ is given by:
$\mathcal{D}_{\mathbf{X}}=\left\{\mathbf{D}=\operatorname{diag}(\mathbbm{1}(\mathbf{Xs}\succeq\mathbf{0})):\mathbf{s}\in\mathbb{R}^{d}\right\},$ (1)
where $\mathbf{D}$ is a diagonal matrix with a possible activation pattern encoded on the diagonal, $\mathbbm{1}$ is the indicator function, and $\succeq$ is an element-wise $\geq$ operator.
By introducing cone-constrained variables $\bm{\nu},\bm{\omega}\in\mathbb{R}^{P\times d}$ and the diagonal matrices from $\mathcal{D}_{\mathbf{X}}$[18] have obtained a convex training procedure from which we can find the previous non-convex weights under some conditions. The optimization problem is now defined as
$p^{\star}=$ $\min_{\bm{\nu},\bm{\omega}}$ $\mathcal{L}\left(\sum_{D_{i}\in\mathcal{D}_{\mathbf{X}}}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i}),\mathbf{y}\right)$ (SCNNT)
@@
where $\mathcal{K}_{i}=\{\mathbf{u}\in\mathbb{R}^{d}:(2\mathbf{D}_{i}-\mathbf{I})\mathbf{X}\mathbf{u}\succeq 0\}$ and $P=\operatorname{card}(\mathcal{D}_{\mathbf{X}})$. For $m\geq m^{\star}=\sum_{i:\bm{\nu}_{i}\neq\mathbf{0}}^{P}1+\sum_{i:,\bm{\omega}_{i}\neq\mathbf{0}}^{P}1$, problems (SNNT) and (SCNNT) have identical optimal values [18]. The relation between the optimal weights of each problem is then given by:
@@
To make the problem computationally tractable, we can relax the full hyperplane arrangements by sampling random vectors [19]. For example, using a Gaussian sampling, we substitute $\mathcal{D}_{\mathbf{X}}$ by $\tilde{\mathcal{D}}_{\mathbf{X}}=\{\mathbf{D}_{i}=\operatorname{diag}(\mathbbm{1}(\mathbf{XS}_{i}\succeq\mathbf{0})):\mathbf{S}_{i}\thicksim\mathcal{N}_{d}(\bm{\mu},\bm{\Sigma}),\mathbf{S}_{i}=\mathbf{s}_{i},\mathbf{D}_{i}\neq\mathbf{D}_{j\neq i},i\in\llbracket P\rrbracket\}$ where $\mathbf{S}_{i}$ is a random Gaussian vector of mean $\bm{\mu}\in\mathbb{R}^{d}$ and of covariance $\bm{\Sigma}\in\mathbb{R}^{d\times d}$, and $\mathbf{s}_{i}$ is the realization of random vector $\mathbf{S}_{i}\ \forall i\in\llbracket P\rrbracket$. This relaxation has the same minima as (SCNNT) if
$m\geq b\coloneq\sum_{\mathbf{D}_{i}\in\tilde{\mathcal{D}}}\{\bm{\nu}_{i}^{\star}:\bm{\nu}_{i}^{\star}\neq\mathbf{0}\}\cup\{\bm{\omega}_{i}^{\star}:\bm{\omega}_{i}^{\star}\neq\mathbf{0}\},$ (4)
where $\bm{\nu}^{\star},\bm{\omega}^{\star}$ are global optima of (SCNNT) and the optimal activations are in the convex model. Interested readers are referred to [18, 19] for the detailed coverage of SCNNs. We remark that if this condition is met, for a feature sample $\mathbf{x}$ not included in the training set $\mathbf{X}$, then the non-convex formulation of the neural network gives the same output as the convex formulation when using the same set of realized sampling vectors $\mathcal{S}=\{\mathbf{S}_{1}=\mathbf{s}_{1},\mathbf{S}_{2}=\mathbf{s}_{s},\ldots,\mathbf{S}_{P}=\mathbf{s}_{P}\}$. This notion will be useful in the next section.
3 WaDiRo SCNNs
We now introduce WaDiRo-SCNNs and provide tractable procedures to train them. First, we reexpress the training function of the SCNN in a form compatible with the tractable formulation of the Wasserstein DRO problem from [25] to obtain performance guarantees for critical applications.

### v2

@@ 1.3 Contributions
In this work, we propose a new distributionally robust convex training procedure for $\operatorname{ReLU}$ shallow neural networks under the Wasserstein metric. Under some conditions, we show that the standard $\operatorname{ReLU}$ shallow convex neural network (SCNN) training can be decoupled and formulated as a constrained linear regression over a modified sample space. SCNN training requires a set of sampling vectors, which are known a priori, to model possible activation patterns. We condition over this set to construct a new independent and identically distributed conditional empirical distribution based on the modified sample space. Decoupling the training and working with this new empirical distribution allows us to easily adapt the SCNN training to the tractable formulation of the Wasserstein-DRO problem proposed by Mohajerin Esfahani and Kuhn (2018), derive out-of-sample performance guarantees stemming from its Rademacher complexity, and enforce hard convex physical constraints during training.
Additionally, we propose a first post-training verification program specifically tailored to SCNNs. We model the inner workings of SCNNs as a mixed-integer convex optimization problem to certify the worst-case local instability in the feature space of a trained SCNN. This post-verification framework enables us to evaluate how different regularization paradigms, such as distributionally robust learning, influence the stability of trained SCNNs without having to map their weights to their non-convex equivalents. This program also opens research opportunities for bilevel programs where SCNN certification and training are tackled simultaneously.
We propose a new methodology to benchmark distributionally robust (and regularized) machine learning models in a controlled synthetic environment. We design a method to introduce real out-of-sample outliers under the sliced-Wasserstein distance (Bonnotte 2013), a computationally tractable approximation of the Wasserstein distance. We combine this method with noise to corrupt machine learning datasets generated with standard optimization benchmark functions typically reserved for non-convex solver benchmarking. We show that our WaDiRo-SCNN training performs well empirically in highly corrupted environments: it is robust and conservative while being able to model non-convex functions. Finally, we test our model on a real-world VPP application, viz., the hourly energy consumption prediction of non-residential buildings from Montréal, Canada. For this application, we enforce the physical constraint that energy consumption is non-negative. We observe that this constraining greatly reduces the number of constraint violations during testing while having no significant precision drop with the original non-physics constrained WaDiRo-SCNN.
To the best of our knowledge, our specific contributions are as follows:
  • We demonstrate that the SCNN with $\ell_{1}$-loss training problem is similar to a constrained linear regression over a modified sample space.
  • We formulate an exact, tractable, and low-stochasticity Wasserstein distributionally robust training program for shallow convex neural networks under both $\ell_{1}$-norm and $\ell_{2}$-norm order-1 Wasserstein distances by exploiting this constrained linear regression formulation and linking it with past literature.
  • We derive out-of-sample guarantees for WaDiRo-SCNNs with $\ell_{1}$-loss training via the Rademacher complexity under its original sample space and its modified one.
@@
  • We design a mixed-integer convex program (MICP) for the post-training verification of SCNNs. We utilize it to show how distributionally robust and regularized training methods increase model stability empirically.
  • We propose a new benchmarking procedure with synthetic data to empirically evaluate the robustness of ML models regarding out-of-distribution outliers.
@@
The rest of the paper is organized as follows. In Section 2, we introduce Wasserstein distributionally robust optimization and its tractable formulation for convex problems (Section 2.1) as well as shallow convex neural networks (Section 2.2). In Section 3 we present the main results of this paper, viz., WaDiRo-SCNNs. In Sections 4, 5, and 6 we address the theoretical out-of-sample performance of our proposition, physics-constrained SCNNs, and SCNN post-training verification, respectively. Finally, in Section 7 we present a methodology to generate outliers and in Section 8, we present all numerical experiments.
@@ 2.1 Distributionally Robust Optimization
Distributionally robust optimization (DRO) minimizes a worst-case expected loss function over an ambiguity set, i.e., an uncertainty set in the space of probability distributions (Kuhn et al. 2019). The problem can be formulated as:
$\inf_{\bm{\beta}}\sup_{\mathbb{Q}\in\Omega}\ \mathbb{E}^{\mathbb{Q}}[h_{\bm{\beta}}(\mathbf{z})],$ (DRO)
where $\mathbf{z}\in\mathcal{Z}\subseteq\mathbb{R}^{d},d\in\mathbb{N}$, is a random sample, $\mathcal{Z}$ is the set of all possible values of $\mathbf{z}$, $\mathbb{Q}$ is a distribution from the ambiguity set $\Omega$, $\bm{\beta}\in\mathbb{R}^{p}$ is the $p$-dimension decision vector, and $h_{\bm{\beta}}(\mathbf{z}):\mathcal{Z}\times\mathbb{R}^{p}\to\mathbb{R}$ is the objective function when applying $\bm{\beta}$ on a sample (Chen and Paschalidis 2020, Mohajerin Esfahani and Kuhn 2018).
We are interested in a data-driven branch of DRO which defines the ambiguity set as a Wasserstein ball, a ball which contains all distributions within a Wasserstein distance from a reference distribution (Yue et al. 2022), centered on an empirical distribution containing $N$ samples: ^$\hat{\mathbb{P}}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathbf{z}_{i}}(\mathbf{z})$, where $\delta_{\mathbf{z}_{i}}(\cdot)$ is a Dirac delta function assigning a probability mass of one at each known sample $\mathbf{z}_{i}$. Note that as the number of samples goes to infinity the empirical distribution converges to the true distribution $\mathbb{P}$ (van der Vaart 2000, Chapter 19.2). The Wasserstein distance, also called earth mover’s distance and closely related to the optimal transport plan (Kuhn et al. 2019), can be pictured as the minimal effort that it would take to displace a given pile of a weighted resource, e.g., dirt, to shape a second specific pile. The effort of a trip is given by the distance traveled and the weight of the carried resource. Specifically, it provides a sense of similarity between the two distributions. Formally, the probabilistic definition of the order-$t$ Wasserstein distance between probability measures $\mathbb{U}$ and $\mathbb{V}$ on $\mathcal{Z}$ is:
$W_{\|\cdot\|,t}(\mathbb{U},\mathbb{V})=\inf_{\begin{subarray}{c}U\sim\mathbb{U},\\ V\sim\mathbb{V}\end{subarray}}\left(\mathbb{E}[\|U-V\|^{t}]\right)^{\frac{1}{t}},\quad t\geq 1,$
where the infimum is taken over all possible couplings of random vectors $U$ and $V$ which are marginally distributed as $\mathbb{U}$ and $\mathbb{V}$, respectively (Panaretos and Zemel 2019). As such, the Wasserstein ball defines a space containing all possible distributions close to the distribution in its center. We refer interested readers to the open-access book of Panaretos and Zemel (2020) which covers the Wasserstein distance thoroughly while offering interesting statistical insights.
@@
where $\Omega$ is the order-$t$ Wasserstein ball of radius $\epsilon>0$ centered on ^$\hat{\mathbb{P}}_{N}$, and $\mathcal{P}(\mathcal{Z})$ is the space of all probability distributions with support on $\mathcal{Z}$. In a way, $\epsilon$ can be interpreted as a hyperparameter that controls our conservatism towards ^$\hat{\mathbb{P}}_{N}$ since a greater radius admits distributions farther from ^$\hat{\mathbb{P}}_{N}$ into the ambiguity set. This problem is not tractable (Mohajerin Esfahani and Kuhn 2018).
Using the strong duality theorem and the order-1 Wasserstein metric, Mohajerin Esfahani and Kuhn (2018) show that an exact tractable reformulation of (WDRO) is obtainable if $h_{\bm{\beta}}(\mathbf{z})$ is convex in $\mathbf{z}\in\mathcal{Z}$ and if $\mathcal{Z}=\mathbb{R}^{d}$, where $d$ is the dimension of the samples. The problem to consider then becomes:
@@
where $\|\cdot\|_{*}$ is the dual norm $\|\bm{\theta}\|_{*}\equiv\sup_{\|\mathbf{z}\|\leq 1}\bm{\theta}^{\top}\mathbf{z}$ and $h_{\bm{\beta}}^{*}(\bm{\theta})$ is the convex conjugate of $h_{\bm{\beta}}(\mathbf{z})$, $h_{\bm{\beta}}^{*}(\bm{\theta})=\sup_{\mathbf{z}}\{\bm{\theta}^{\top}\mathbf{z}-h_{\bm{\beta}}(\mathbf{z})\}$ (Boyd and Vandenberghe 2004).
Training machine learning models with (TWDRO) has been proved to be a generalization of regularization (Shafieezadeh-Abadeh et al. 2019). By training over the worst-case expected distribution in a Wasserstein ball centered on the training dataset, we can expect to avoid overfitting on the empirical distribution and yielding increased robustness against outliers, noise, and some adversarial attacks.
@@ 2.2 Shallow Convex Neural Networks
Let $\mathbf{X}\in\mathbb{R}^{N\times d}$ be a data matrix of $N$ feature vectors $\mathbf{x}_{j}\in\mathbb{R}^{d}\ \forall j\in\llbracket N\rrbracket$, and $\mathbf{y}\in\mathbb{R}^{N}$ be the corresponding one-dimensional labels. Define data samples as a feature-label pair: $\mathbf{z}_{j}=(\mathbf{x}_{j},y_{j})\ \forall j\in\llbracket N\rrbracket$. The non-convex training problem of a single-output feedforward shallow neural network (SNN) with $\operatorname{ReLU}$ activation functions is as follows:
$p^{\star}$ $=\min_{\mathbf{W}_{1},\mathbf{w}_{2}}\mathcal{L}\left(\sum_{i=1}^{m}\{(\mathbf{X}\mathbf{W}_{1i})_{+}w_{2i}\},\mathbf{y}\right),$ (SNNT)
where $m$ is the number of hidden neurons, $\mathbf{W}_{1}\in\mathbb{R}^{d\times m}$, and $\mathbf{w}_{2}\in\mathbb{R}^{m}$ are the weights of the first and second layers, $\mathcal{L}:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is a convex loss function, and $(\cdot)_{+}$ is an element-wise $\max\{0,\cdot\}$ operator ($\operatorname{ReLU}$).
The convex reformulation of the SNN training problem (SNNT) is based on an enumeration process of all the possible $\operatorname{ReLU}$ activation patterns in the hidden layer for a fixed training dataset $\mathbf{X}$ (Pilanci and Ergen 2020). The enumeration is equivalent to generating all possible hyperplane arrangements passing through the origin and clustering the data with them. The set of all possible activation patterns a $\operatorname{ReLU}$ SNN can take for a training set $\mathbf{X}$ is given by:
$\mathcal{D}_{\mathbf{X}}=\left\{\mathbf{D}=\operatorname{diag}(\text{1}(\mathbf{Xs}\succeq\mathbf{0})):\mathbf{s}\in\mathbb{R}^{d}\right\},$ (1)
where $\mathbf{D}$ is a diagonal matrix with a possible activation pattern encoded on the diagonal.
By introducing the cone-constrained variables $\bm{\nu},\bm{\omega}\in\mathbb{R}^{P\times d}$ and the diagonal matrices from $\mathcal{D}_{\mathbf{X}}$ilanci and Ergen (2020) have obtained a convex training procedure from which we can retrieve the previous non-convex weights under some conditions. The optimization problem is now defined as
$p^{\star}=$ $\min_{\bm{\nu},\bm{\omega}}$ $\mathcal{L}\left(\sum_{i=1}^{P}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i}),\mathbf{y}\right)$ (SCNNT)
@@
where $\mathcal{K}_{i}=\{\mathbf{u}\in\mathbb{R}^{d}:(2\mathbf{D}_{i}-\mathbf{I})\mathbf{X}\mathbf{u}\succeq\mathbf{0}\}$ and $P=\operatorname{card}(\mathcal{D}_{\mathbf{X}})$. For $m\geq m^{\star}=\sum_{i:\bm{\nu}_{i}\neq\mathbf{0}}^{P}1+\sum_{i:,\bm{\omega}_{i}\neq\mathbf{0}}^{P}1$, with $11\leq m^{\star}\leq N+$ 1, problems (SNNT) and (SCNNT) have identical optimal values (Pilanci and Ergen 2020). At least one optimal non-convex NN from (SNNT) can be assembled, using $m^{\star}$ neurons, with the optimal solution of (SCNNT) and the following mapping between the optimal weights of each problem:
@@
We note that Wang et al. (2021) demonstrate how to recover the full optimal set of non-convex networks.
To make the problem computationally tractable, we can approximate the full hyperplane arrangements by sampling random vectors (Mishkin et al. 2022). For example, using a Gaussian sampling, we substitute $\mathcal{D}_{\mathbf{X}}$ by $\tilde{\mathcal{D}}_{\mathbf{X}}=\{\mathbf{D}_{i}=\operatorname{diag}(\text{1}(\mathbf{XS}_{i}\succeq\mathbf{0})):\mathbf{S}_{i}\thicksim\mathcal{N}_{d}(\bm{\mu},\bm{\Sigma}),\mathbf{D}_{i}\neq\mathbf{D}_{j\neq i},i\in\llbracket P\rrbracket\}$ where $\mathbf{S}_{i}$ is a random Gaussian vector of mean $\bm{\mu}\in\mathbb{R}^{d}$ and of covariance $\bm{\Sigma}\in\mathbb{R}^{d\times d}$, and such that $\mathbf{s}_{i}\in\mathbb{R}^{d}$ is its realization, $\forall i\in\llbracket P\rrbracket$. This approximation has the same minima as (SCNNT) if
$m\geq b\coloneq\sum_{\mathbf{D}_{i}\in\tilde{\mathcal{D}}_{\mathbf{X}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\operatorname{card}}}\left(\{\bm{\nu}_{i}^{\star}:\bm{\nu}_{i}^{\star}\neq\mathbf{0}\}\cup\{\bm{\omega}_{i}^{\star}:\bm{\omega}_{i}^{\star}\neq\mathbf{0}\}\right),$
where in this case $\bm{\nu}^{\star},\bm{\omega}^{\star}$ are global optima of the sub-sampled convex problem and if the optimal activations are in the convex model, i.e., $\left\{\operatorname{diag}(\mathbf{XW}_{1i}^{\star}\geq 0):i\in\llbracket m\rrbracket\right\}\subseteq\tilde{\mathcal{D}}_{X}$ (Mishkin et al. 2022). Interested readers are referred to Pilanci and Ergen (2020) and Mishkin et al. (2022) for the detailed coverage of SCNNs.
We remark that if these conditions are met, for a feature sample $\mathbf{x}$ not included in the training dataset $\mathbf{X}$, then the non-convex formulation of the neural network gives the same output as the convex formulation when using the same set of realized sampling vectors $\mathcal{S}=\{\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{P}\}$. This notion will be useful in the next section.
If these conditions are not met, i.e., the SNN is not wide enough, we still obtain an approximately optimal network that can outperform conventional training methods as highlighted by Mishkin et al. (2022). By approximately optimal, we mean that it is optimized by a convex program that guarantees optimality while not having an exact mapping to the optimal non-convex neural network formulation. Thus, choosing the maximal width of the SCNN is a way to control the tractability of the convex training with respect to the available computational power and the size of the training dataset.
3 Wasserstein Distributionally Robust Shallow Convex Neural Networks
We now introduce WaDiRo-SCNNs and provide tractable procedures to train them. First, we reexpress the training objective function of the SCNN in a form compatible with the tractable formulation of the Wasserstein DRO problem from Mohajerin Esfahani and Kuhn (2018) to obtain performance guarantees for critical applications.